# Conformal four-point ladder integrals in diverse dimensions and polylogarithms

### v1

@@ Introduction
$\begin{array}[]{l}\displaystyle\Phi^{(\beta)}_{L}({\sf x},{\sf y})=\dfrac{\Gamma(\lambda)\,L!\,4^{L}{\sf x}^{2\beta}}{2\pi^{\lambda+2}({\sf x}^{2}\,{\sf y}^{2})^{D/4}}\,\sum\limits_{n=0}^{\infty}(n+\lambda)\,C^{(\lambda)}_{n}\left(\hat{{\sf x}}\hat{{\sf y}}\right)\,\;\times\\[8.5359pt] \displaystyle\qquad\qquad\qquad\qquad\qquad\qquad\times\;\int\limits_{-\infty}^{+\infty}d\nu\,\frac{({\sf y}^{2}/{\sf x}^{2})^{{\sf i}\nu}}{\Bigl(4^{\beta}\prod\limits_{m=1}^{\beta}(\frac{D}{4}+\frac{n}{2}+{\sf i}\nu-m)(\frac{D}{4}+\frac{n}{2}-1+m-{\sf i}\nu)\Bigr)^{L+1}}\;,\end{array}$ (3.1)
where $\lambda=D/2-1$. Further, in this subsection we consider the simplest case $\beta=\alpha=1$ ($\gamma=D/2-1$). This choice of parameters in the case $D=4$ leads to the usual ladder diagrams with all indices on the lines equal to $1$. For arbitrary dimension $D$ the equation (3.1) represents the ladder diagram in Fig. 1 with indices $1$ on the horizontal lines and D/21D/2-1 on the vertical lines. The corresponding integrals are convergent for any $D$ and we deduce from eq. (3.1) the expression
@@
Integrating over $\nu$ we obtain the form [^3] We close the contour of integration in the lower half-plane, assuming ${\sf x}^{2}>{\sf y}^{2}$. The case of ${\sf x}^{2}<{\sf y}^{2}$ is automatically taken into account due to the symmetry (2.17). (see eq. (3.47) in [70] [^4] In [70] we used notation $u,w$ for vectors ${\sf x},{\sf y}$.)
@@
is a Lerch function (see [89]), which generalizes the polylogarithms
@@
Proof. First, we rewrite our expression (3.5) in the following form
@@
where we apply an obvious identity $\left.f(x\partial_{x})\,x^{n}\right|_{x=1}=f(n)$ and substitute the explicit expression (2.37) of generating function of Gegenbauer polynomials, for $t=x\sqrt{z\bar{z}}$ and $r=\frac{z+\bar{z}}{2\sqrt{z\bar{z}}}$. Then, we note that
@@
where operator $P_{\lambda}(z\partial_{z})$ was introduced in (3.7). Indeed, we have
@@
We use the definition of Lerch function (3.8):
@@
Note that the operator of dilatation $z\partial_{z}$ applied to the Lerch function shifts its parameter, namely
@@
Thus, application of the differential operator in (3.6) can be reduced to the simple transformations of Lerch functions $\Phi(z,s,\lambda)$. In what follows we will limit ourselves with the case of positive integer $\lambda=D/2-1\in\mathbb{Z}_{>0}$, which corresponds to the even dimensions $D=2(1+\lambda)$. For such a choice of $\lambda$, the operator $P_{\lambda}(z\partial_{z})$ in (3.16) becomes a polynomial in $z\partial_{z}$:
@@
with $P_{1}(z\partial_{z})=1$ and one can directly apply formula (3.15) in (3.6), (3.14) to obtain an explicit expression for $\Sigma_{s}^{(\lambda)}(z,\bar{z})$. We note that the representation in the right hand side of (3.16), which follows form the fact $\big[\partial_{z},z^{p}\big]=pz^{p-1}$, is more convenient for the analytical continuation of $P_{\lambda}(z\partial_{z})$ for $\lambda\in\mathbb{C}$.
@@
In the case $\lambda=D/2-1\in\mathbb{Z}_{>0}$, the function $\Sigma_{s}^{(\lambda)}(z,\bar{z})$ defined in (3.6) has explicit expression via polylogarithms
@@
where operators $P_{\lambda}(z\partial_{z})$ and $P_{\lambda}(\bar{z}\partial_{\bar{z}})$ were defined in (3.16).
Proof. For $\lambda\in\mathbb{Z}_{>0}$, Lerch function (3.8) is reduced to the sum of polylogarithm and polynomial in $z^{-1}$. Indeed, for $\lambda>1$ we have
@@
Applying this relation to (3.14) we get
@@
Here we use the identity [^5] For $\lambda=1,2$ the expression in the left hand-side of (3.20) automatically vanishes. $(\lambda>2)$
@@
which follows from the fact that the operator $P_{\lambda}(z\partial_{z})$ defined in (3.16) and its analog $P_{\lambda}(\bar{z}\partial_{\bar{z}})$ have zero modes $z^{-1}\,,z^{-2}\,,...\,,z^{1-\lambda}$ and $\bar{z}^{-1}\,,\bar{z}^{-2}\,,...\,,\bar{z}^{1-\lambda}$. Thus, the sum over $m$ appeared in (3.20) evidently belongs to the kernel of the product of these operators.
Finally we apply formula (3.10) and its analog for $z\to\bar{z}$ to equation (3.19) to obtain (3.17). ∎
We note that the substitution of (3.19) into (3.4) reproduces (up to a normalization) the result [88, eq. (3.6)].
@@
The functions (3.17) satisfy relation
@@
where recursion operator $R_{d}$ was defined in (2.35).
@@
and in view of the symmetry $R_{d}|_{z\leftrightarrow\bar{z}}=R_{d}$ we need only to prove identity
@@
where in the last equality use identity
@@
which follows from obvious relation $\partial_{z}^{\lambda}(z-\bar{z})=(z-\bar{z})\partial_{z}^{\lambda}+\lambda\partial_{z}^{\lambda-1}$.
We argue that representation of the function $\Sigma_{s}^{(\lambda)}(z,\bar{z})$ in the form (3.17) is already enough to build the recurrent procedure for obtaining explicit expressions for any $\lambda\in\mathbb{Z}_{\geq 0}$. In order to proceed with such calculations it might be useful to note the property of the dilatation operator $z\partial_{z}$, its application to the polylogarithm shifts its weight
@@
Remark 8. Note that despite the relation (2.36) and (3.21) were derived from the different starting points, namely, one from the general representation (2.31) for generic $\beta$, and another from the application of differential operator $P_{\lambda}(z\partial_{z})$ in (3.17), we can easily see their connection. Operator $R_{d}$ commutes with any function $f(z\bar{z})$ and in particular with the $\log(z\bar{z})$,
@@
that is needed to shift the application of this operator from $\Sigma_{s}^{(\lambda)}(z,\bar{z})$ to the whole function $\tilde{\Phi}_{L}^{(\lambda)}(z,\bar{z};\lambda)$.
Remark 9. In view of relation (3.4) the function $\Phi_{L}({\sf x},{\sf y})$ depends only on the product $\widetilde{\Sigma}_{s}^{(\lambda)}=\Gamma(\lambda)\Sigma_{s}^{(\lambda)}$. According to (3.21) the function $\widetilde{\Sigma}_{s}^{(\lambda)}$ satisfies equation
@@
Using this relation we can define the two-dimensional function
@@
where $X(z\bar{z})$ is the zero mode of the operator $R_{d}$. One can see that the two-dimensional function (2.33) is singular in the case $\beta=1$, which means that in order to relate (2.33) and (3.25) $X(z\bar{z})$ should correspond to singular contribution. For more details see subsection 3.3.
@@
    $\Sigma^{(2)}_{s}(z,\bar{z})=(z\partial_{z}+1)\dfrac{\operatorname{Li}_{s}(z)}{(z-\bar{z})^{2}}+(z\leftrightarrow\bar{z})=-\dfrac{z+\bar{z}}{(z-\bar{z})^{3}}\operatorname{Li}_{s}(z)+\dfrac{\operatorname{Li}_{s-1}(z)}{(z-\bar{z})^{2}}+(z\leftrightarrow\bar{z}).$
@@
    $\Sigma_{s}^{(3)}(z,\bar{z})=\dfrac{1}{2}(z\partial_{z}+2)(z\partial_{z}+1)\dfrac{\operatorname{Li}_{s}(z)}{(z-\bar{z})^{3}}+(z\leftrightarrow\bar{z})\\ =\dfrac{z^{2}+4z\bar{z}+\bar{z}^{2}}{(z-\bar{z})^{5}}\operatorname{Li}_{s}(z)-\dfrac{3(z+\bar{z})}{2(z-\bar{z})^{4}}\operatorname{Li}_{s-1}(z)+\dfrac{\operatorname{Li}_{s-2}(z)}{2(z-\bar{z})^{3}}+(z\leftrightarrow\bar{z}).$
@@
    $\Sigma_{s}^{(4)}(z,\bar{z})=\dfrac{1}{6}(z\partial_{z}+3)(z\partial_{z}+2)(z\partial_{z}+1)\dfrac{\operatorname{Li}_{s}(z)}{(z-\bar{z})^{4}}+(z\leftrightarrow\bar{z})\\ =-\dfrac{(z+\bar{z})(z^{2}+8z\bar{z}+\bar{z}^{2})}{(z-\bar{z})^{7}}\operatorname{Li}_{s}(z)+\dfrac{11z^{2}+38z\bar{z}+11\bar{z}^{2}}{6(z-\bar{z})^{6}}\operatorname{Li}_{s-1}(z)\\ -\dfrac{z+\bar{z}}{(z-\bar{z})^{5}}\operatorname{Li}_{s-2}(z)+\dfrac{\operatorname{Li}_{s-3}(z)}{6(z-\bar{z})^{4}}+(z\leftrightarrow\bar{z}).$
@@
    $\Sigma_{s}^{(5)}(z,\bar{z})=\frac{1}{24}(z\partial_{z}+4)(z\partial_{z}+3)(z\partial_{z}+2)(z\partial_{z}+1)\frac{1}{(z-\bar{z})^{5}}\,\operatorname{Li}_{s}(z)+(z\leftrightarrow\bar{z})\\ =\frac{z^{4}+16z^{3}\bar{z}+36z^{2}\bar{z}^{2}+16z\bar{z}^{3}+\bar{z}^{4}}{(z-\bar{z})^{9}}\,\operatorname{Li}_{s}(z)-\frac{5(z+\bar{z})(5z^{2}+32z\bar{z}+5\bar{z}^{2})}{12(z-\bar{z})^{8}}\,\operatorname{Li}_{s-1}(z)\\ +\frac{5(7z^{2}+22z\bar{z}+7\bar{z}^{2})}{24(z-\bar{z})^{7}}\,\operatorname{Li}_{s-2}(z)-\frac{5(z+\bar{z})}{12(z-\bar{z})^{6}}\operatorname{Li}_{s-3}(z)+\dfrac{\operatorname{Li}_{s-4}(z)}{24(z-\bar{z})^{5}}+(z\leftrightarrow\bar{z}).$
Analyzing examples for the small values of $\lambda$ we can clearly see the pattern. For a given $\lambda$, answer for $\Sigma^{(\lambda)}_{s}$ is expressed via number of polylogarithms $\operatorname{Li}_{s}(z),\operatorname{Li}_{s-1}(z),\ldots,\operatorname{Li}_{s-\lambda+1}(z)$ accompanied with rational functions of the form $\frac{G_{k}(z,\bar{z})}{(z-z)^{2\lambda-k-1}}$, where $G_{k}(z,\bar{z})$ is homogeneus polynomial of degree $\lambda-1-k$ symmetrical under the transformation $z\leftrightarrow\bar{z}$. In Appendix C we give a compact formula for the polynomial $G_{k}^{(\lambda)}(z,\bar{z})$ and show that such properties as homogeneousness and symmetry holds for any $\lambda\in\mathbb{Z}_{>0}$.
At the end of this subsection we give explicit expressions for the function (2.31) for $\beta=1$ in $D=4,6,8,10$.
1. The case $D=4$ ($\lambda=D/2-1=1$) and indices on the lines $\alpha=\beta=\gamma=1$ [60, 61] (see also [58, 70]):
@@
2. The case $D=6$ ($\lambda=2$) and indices on the lines $\alpha=\beta=1$, $\gamma=2$ [76]:
$\begin{array}[]{l}\widetilde{\Phi}^{(1)}_{L}(z,\bar{z};2)=\dfrac{\pi\,(z\bar{z})^{1/2}}{2(z-\bar{z})^{2}\,4^{L}}\sum\limits_{k=0}^{L}C^{L}_{2L-k}\,\dfrac{(-1)^{k}\log^{k}(z\bar{z})}{k!}\\[8.5359pt] \qquad\qquad\Bigl(-\dfrac{(z+\bar{z})}{(z-\bar{z})}(\operatorname{Li}_{2L-k}(z)-\operatorname{Li}_{2L-k}(\bar{z}))+\operatorname{Li}_{2L-k-1}(z)+\operatorname{Li}_{2L-k-1}(\bar{z})\Bigr)\,.\end{array}$ (3.27)

### v2

@@ 3.1 Ladder integrals and polylogarithms for $\beta=1$
$\begin{array}[]{l}\displaystyle\Phi^{(\beta)}_{L}({\sf x},{\sf y})=\dfrac{\Gamma(\lambda)\,L!\,4^{L}{\sf x}^{2\beta}}{2\pi^{\lambda+2}({\sf x}^{2}\,{\sf y}^{2})^{D/4}}\,\sum\limits_{n=0}^{\infty}(n+\lambda)\,C^{(\lambda)}_{n}\left(\hat{{\sf x}}\hat{{\sf y}}\right)\,\;\times\\[8.5359pt] \displaystyle\quad\quad\quad\quad\quad\quad\times\;\int\limits_{-\infty}^{+\infty}d\nu\,\frac{({\sf y}^{2}/{\sf x}^{2})^{{\sf i}\nu}}{\Bigl(4^{\beta}\prod\limits_{m=1}^{\beta}(\frac{D}{4}+\frac{n}{2}+{\sf i}\nu-m)(\frac{D}{4}+\frac{n}{2}-1+m-{\sf i}\nu)\Bigr)^{L+1}}\;,\end{array}$ (3.1)
where $\lambda=D/2-1$. Further, in this subsection, we consider the simplest case $\beta=\alpha=1$ ($\gamma=D/2-1$). This choice of parameters in the case $D=4$ leads to usual ladder diagrams with all indices on the lines equal to $1$. For arbitrary dimension $D$ the equation (3.1) represents the ladder diagram in Fig. 1 with indices $1$ on the horizontal lines and D/21D/2-1 on the vertical lines. The corresponding integrals are convergent for any $D$, and we deduce from eq. (3.1) the expression
@@
Integrating over $\nu$, we obtain the form§§§We close the contour of integration in the lower half-plane, assuming ${\sf x}^{2}>{\sf y}^{2}$. The case of ${\sf x}^{2}<{\sf y}^{2}$ is automatically taken into account due to the symmetry (2.17). (see eq. (3.47) in Derkachov et al. (2023)In Derkachov et al. (2023), we used the notation $u,w$ for the vectors ${\sf x},{\sf y}$.)
@@
is the Lerch function (see Erdélyi and Bateman (1953a)), which generalizes the polylogarithms
@@
Proof. First, we rewrite our expression (3.5) in the following form:
@@
where we apply an obvious identity $\left.f(x\partial_{x})\,x^{n}\right|_{x=1}=f(n)$ and substitute the explicit expression (2.37) of the generating function of Gegenbauer polynomials for $t=x\sqrt{z\bar{z}}$ and $r=\frac{z+\bar{z}}{2\sqrt{z\bar{z}}}$. Then, we note that
@@
where the operator $P_{\lambda}(z\partial_{z})$ was introduced in (3.7). Indeed, we have
@@
We use the definition of the Lerch function (3.8):
@@
Note that the operator of dilatation $z\partial_{z}$ applied to the Lerch function shifts its parameter, namely,
@@
Thus, the application of the differential operator in (3.6) can be reduced to simple transformations of the Lerch functions $\Phi(z,s,\lambda)$. In what follows we will limit ourselves to the case of positive integer $\lambda=D/2-1\in\mathbb{Z}_{>0}$, which corresponds to the even dimensions $D=2(1+\lambda)$. For such a choice of $\lambda$, the operator $P_{\lambda}(z\partial_{z})$ in (3.16) becomes a polynomial in $z\partial_{z}$:
@@
with $P_{1}(z\partial_{z})=1$ and one can directly apply formula (3.15) in (3.6), (3.14) to obtain an explicit expression for $\Sigma_{s}^{(\lambda)}(z,\bar{z})$. We note that the representation in the right-hand side of (3.16), which follows form the fact $\big[\partial_{z},z^{p}\big]=pz^{p-1}$, is more convenient for the analytical continuation of $P_{\lambda}(z\partial_{z})$ for $\lambda\in\mathbb{C}$.
@@
In the case $\lambda=D/2-1\in\mathbb{Z}_{>0}$, the function $\Sigma_{s}^{(\lambda)}(z,\bar{z})$ defined in (3.6) has an explicit expression in terms of polylogarithms
@@
where the operators $P_{\lambda}(z\partial_{z})$ and $P_{\lambda}(\bar{z}\partial_{\bar{z}})$ were defined in (3.16).
Proof. For $\lambda\in\mathbb{Z}_{>0}$, the Lerch function (3.8) is reduced to the sum of a polylogarithm and a polynomial in $z^{-1}$. Indeed, for $\lambda>1$ we have
@@
Applying this relation to (3.14), we get
@@
Here we use the identityFor $\lambda=1,2$ the expression in the left hand-side of (3.20) automatically vanishes. $(\lambda>2)$
@@
which follows from the fact that the operator $P_{\lambda}(z\partial_{z})$ defined in (3.16) and its analog $P_{\lambda}(\bar{z}\partial_{\bar{z}})$ have zero modes $z^{-1}\,,z^{-2}\,,...\,,z^{1-\lambda}$ and $\bar{z}^{-1}\,,\bar{z}^{-2}\,,...\,,\bar{z}^{1-\lambda}$. Thus, the sum over $m$ appearing in (3.20) evidently belongs to the kernel of the product of these operators.
Finally, we apply formula (3.10) and its analog for $z\to\bar{z}$ to equation (3.19) to obtain (3.17). ∎
We note that the substitution of (3.19) into (3.4) reproduces (up to normalization) the result (Borinsky and Schnetz, 2022b, eq. (3.6)).
@@
The functions (3.17) satisfy the relation
@@
where the recursion operator $R_{d}$ was defined in (2.35).
@@
and in view of the symmetry $R_{d}|_{z\leftrightarrow\bar{z}}=R_{d}$, we need only to prove the identity
@@
In the last equality we use the identity
@@
which follows from the obvious relation $\partial_{z}^{\lambda}(z-\bar{z})=(z-\bar{z})\partial_{z}^{\lambda}+\lambda\partial_{z}^{\lambda-1}$.
We argue that the representation of the function $\Sigma_{s}^{(\lambda)}(z,\bar{z})$ in the form (3.17) is already sufficient to build a recurrent procedure for obtaining explicit expressions for any $\lambda\in\mathbb{Z}_{\geq 0}$. In order to proceed with such calculations, it might be useful to note the property of the dilatation operator $z\partial_{z}$; its application to the polylogarithm shifts its weight
@@
Remark 8. Note that despite the fact that relations (2.36) and (3.21) were derived from the different starting points, namely, one from the general representation (2.31) for generic $\beta$ and the other from the application of the differential operator $P_{\lambda}(z\partial_{z})$ in (3.17), we can easily see their connection. THe operator $R_{d}$ commutes with any function $f(z\bar{z})$ and in particular with $\log(z\bar{z})$,
@@
which is needed to shift the application of this operator from $\Sigma_{s}^{(\lambda)}(z,\bar{z})$ to the whole function $\tilde{\Phi}_{L}^{(\lambda)}(z,\bar{z};\lambda)$.
Remark 9. In view of relation (3.4), the function $\Phi_{L}({\sf x},{\sf y})$ depends only on the product $\widetilde{\Sigma}_{s}^{(\lambda)}=\Gamma(\lambda)\Sigma_{s}^{(\lambda)}$. According to (3.21), the function $\widetilde{\Sigma}_{s}^{(\lambda)}$ satisfies the equation
@@
Using this relation, we can define the two-dimensional function
@@
where $X(z\bar{z})$ is the zero mode of the operator $R_{d}$. One can see that the two-dimensional function (2.33) is singular in the case $\beta=1$, which means that in order to relate (2.33) and (3.25) $X(z\bar{z})$ should correspond to a singular contribution. For more details, see subsection 3.3.
@@
    $\Sigma^{(2)}_{s}(z,\bar{z})=$ $(z\partial_{z}+1)\dfrac{\operatorname{Li}_{s}(z)}{(z-\bar{z})^{2}}+(z\leftrightarrow\bar{z})$
    $=$ $-\dfrac{z+\bar{z}}{(z-\bar{z})^{3}}\operatorname{Li}_{s}(z)+\dfrac{\operatorname{Li}_{s-1}(z)}{(z-\bar{z})^{2}}+(z\leftrightarrow\bar{z}).$
@@
    $\Sigma_{s}^{(3)}(z,\bar{z})=$ $\dfrac{1}{2}(z\partial_{z}+2)(z\partial_{z}+1)\dfrac{\operatorname{Li}_{s}(z)}{(z-\bar{z})^{3}}+(z\leftrightarrow\bar{z})$
    $=\dfrac{z^{2}+4z\bar{z}+\bar{z}^{2}}{(z-\bar{z})^{5}}\operatorname{Li}_{s}(z)-\dfrac{3(z+\bar{z})}{2(z-\bar{z})^{4}}\operatorname{Li}_{s-1}(z)+\dfrac{\operatorname{Li}_{s-2}(z)}{2(z-\bar{z})^{3}}+(z\leftrightarrow\bar{z}).$
@@
    $\Sigma_{s}^{(4)}(z,\bar{z})=$ $\dfrac{1}{6}(z\partial_{z}+3)(z\partial_{z}+2)(z\partial_{z}+1)\dfrac{\operatorname{Li}_{s}(z)}{(z-\bar{z})^{4}}+(z\leftrightarrow\bar{z})$
    $=$ $-\dfrac{(z+\bar{z})(z^{2}+8z\bar{z}+\bar{z}^{2})}{(z-\bar{z})^{7}}\operatorname{Li}_{s}(z)+\dfrac{11z^{2}+38z\bar{z}+11\bar{z}^{2}}{6(z-\bar{z})^{6}}\operatorname{Li}_{s-1}(z)$
    $-\dfrac{z+\bar{z}}{(z-\bar{z})^{5}}\operatorname{Li}_{s-2}(z)+\dfrac{\operatorname{Li}_{s-3}(z)}{6(z-\bar{z})^{4}}+(z\leftrightarrow\bar{z}).$
@@
    $\Sigma_{s}^{(5)}(z,\bar{z})=$ $\frac{1}{24}(z\partial_{z}+4)(z\partial_{z}+3)(z\partial_{z}+2)(z\partial_{z}+1)\frac{1}{(z-\bar{z})^{5}}\,\operatorname{Li}_{s}(z)+(z\leftrightarrow\bar{z})$
    $=$ $\frac{z^{4}+16z^{3}\bar{z}+36z^{2}\bar{z}^{2}+16z\bar{z}^{3}+\bar{z}^{4}}{(z-\bar{z})^{9}}\,\operatorname{Li}_{s}(z)-\frac{5(z+\bar{z})(5z^{2}+32z\bar{z}+5\bar{z}^{2})}{12(z-\bar{z})^{8}}\,\operatorname{Li}_{s-1}(z)$
    $+\frac{5(7z^{2}+22z\bar{z}+7\bar{z}^{2})}{24(z-\bar{z})^{7}}\,\operatorname{Li}_{s-2}(z)-\frac{5(z+\bar{z})}{12(z-\bar{z})^{6}}\operatorname{Li}_{s-3}(z)+\dfrac{\operatorname{Li}_{s-4}(z)}{24(z-\bar{z})^{5}}+(z\leftrightarrow\bar{z}).$
Analyzing the examples for small values of $\lambda$, we can clearly see the pattern. For a given $\lambda$, the answer for $\Sigma^{(\lambda)}_{s}$ is expressed in terms of the number of polylogarithms $\operatorname{Li}_{s}(z),\operatorname{Li}_{s-1}(z),\ldots,\operatorname{Li}_{s-\lambda+1}(z)$ accompanied by rational functions of the form $\frac{G_{k}(z,\bar{z})}{(z-z)^{2\lambda-k-1}}$, where $G_{k}(z,\bar{z})$ is a homogeneus polynomial of degree $\lambda-1-k$ symmetric under the transformation $z\leftrightarrow\bar{z}$. In Appendix C, we give a compact formula for the polynomial $G_{k}^{(\lambda)}(z,\bar{z})$ and show that properties such as homogeneity and symmetry hold for any $\lambda\in\mathbb{Z}_{>0}$.
[…]
1. The case $D=4$ ($\lambda=D/2-1=1$) and indices on the lines $\alpha=\beta=\gamma=1$ Usyukina and Davydychev (1993a, b) (see also Isaev (2003); Derkachov et al. (2023)):
@@
2. The case $D=6$ ($\lambda=2$) and indices on the lines $\alpha=\beta=1$, $\gamma=2$ Loebbert and Stawinski (2024):
$\begin{array}[]{l}\widetilde{\Phi}^{(1)}_{L}(z,\bar{z};2)=\dfrac{\pi\,(z\bar{z})^{1/2}}{2(z-\bar{z})^{2}\,4^{L}}\sum\limits_{k=0}^{L}C^{L}_{2L-k}\,\dfrac{(-1)^{k}\log^{k}(z\bar{z})}{k!}\\[8.5359pt] \quad\quad\Bigl(-\dfrac{(z+\bar{z})}{(z-\bar{z})}(\operatorname{Li}_{2L-k}(z)-\operatorname{Li}_{2L-k}(\bar{z}))+\operatorname{Li}_{2L-k-1}(z)+\operatorname{Li}_{2L-k-1}(\bar{z})\Bigr)\,.\end{array}$ (3.27)